# An Effective Version of the $p$-Curvature Conjecture for Order One Differential Equations

### v1

@@
An Effective Version of the $p$-Curvature Conjecture
for Order One Differential Equations
Florian Fьrnsinn and Lucas Pannier
Аннотация
@@ 1 Introduction
Notably, in his paper from 1980 that popularized the study of D-finite series, i.e., power series that satisfiy a non-zero differential equation of the form (1), Stanley asked for an ‘‘algorithm suitable for computer implementation’’ for problem (P) [Sta80].
Problem (A) was solved algorithmically in a seminal paper by Singer [Sin79], relying on an algorithmic answer of Risch [Ris70] in the case of first order differential equations with algebraic power series coefficients instead of polynomials – this instance of the problem is known as Abel’s problem – and on work conducted by Painlevй and Boulanger [Pai87, Bou97]. However, this answer is not fully satisfactory, as the complexity of the algorithm involves exponential bounds, making it in general not suitable for implementation. Recent progress on Stanley’s question about problem (P) was made by Bostan, Salvy and Singer [BSS25], giving (semi-) algorithms that rely on the minimization of differential operators annihilating a given D-finite series.
A different – arithmetic – approach to problem (A) was proposed by Grothendieck, in his $p$-curvature conjecture [Kat72]. Given (1), we can pass to a matrix differential equation $Y^{\prime}(x)=A(x)Y(x)$ of order $1$, with $A(x)\in\mathbb{Q}(x)^{n\times n}$ being the companion matrix of (1). Let $p$ be a prime number. The $p$-curvature of the equation is defined as the $\mathbb{F}_{p}(x)$-linear map $\psi_{A}:\mathbb{F}_{p}(x)^{n}\to\mathbb{F}_{p}(x)^{n},Y(x)\mapsto(\partial-A_{p})^{p}Y(X)$, where $A_{p}(x)\in\mathbb{F}_{p}(x)^{n\times n}$ denotes the reduction of $A(x)$ modulo $p$. The $p$-curvature conjecture then states that (1) admits a basis of algebraic solutions if and only if its $p$-curvature vanishes for almost all, i.e., all but finitely many, prime numbers $p$. Cartier’s Lemma (as attributed by Katz [Kat72]), asserts that the vanishing of the $p$-curvature is equivalent to the existence of a full basis of (algebraic) solutions of the reduction of the differential equation modulo $p$ in $\mathbb{F}_{p}[\![x]\!]$. Thus, the $p$-curvature conjecture can be seen as a local-global principle about the existence of algebraic solutions of a linear differential equation.
While in general wide open, the $p$-curvature conjecture is solved in many cases. Most notably, for ‘‘equations coming from geometry’’, i.e., suitable factors of Picard-Fuchs differential operators, a proof was given by Katz [Kat72], and for hypergeometric differential equations an elementary proof is a by-product of the classification of algebraic hypergeometric functions [BH89, FY24]. For first order equations
@@
with $u(x)\in\mathbb{Q}(x)$ the three problems (A), (E) and (P) are equivalent, as there is only a one dimensional solution space of the equation. Honda [Hon81] provided an elementary proof of the $p$-curvature conjecture in this case, by showing that the problem is equivalent to a number theoretic result by Kronecker [Kro80], which nowadays is often seen as a consequence of Chebotarev’s Density Theorem [Чеб23a, Чеб23, Tsc26].
@@
For first order equations with algebraic power series coefficients, i.e., for the equation $y^{\prime}(x)=u(x)y(x)$ with $u(x)\in\overline{\mathbb{Q}}[\![x]\!]\cap\overline{\mathbb{Q}(x)}$, a proof of the $p$-curvature conjecture was given by the Chudnovsky brothers [CC85]. They use Hermite-Padй approximation to obtain their result. As a motivating example, they apply their methods to provide a new proof of Kronecker’s Theorem, not relying on Chebotarev’s Density Theorem, and in doing so giving also a new proof of Honda’s Theorem.
The $p$-curvature conjecture, as stated here, does not provide a way to decide whether (1) has a basis of algebraic solutions, as it translates the statement to a statement about almost all, in particular, a statement about infinitely many, prime numbers. In 1982 Katz extended the $p$-curvature conjecture to a conjecture about the Lie algebra of the differential Galois group of a differential equation [Kat82]. From his work presented in [Kat82, §9] it follows by a noetherianity argument that - in the cases in which his conjecture is true, among them order one equations with rational function coefficients - there exists a finite set of prime numbers, such that the vanishing of the $p$-curvatures modulo these primes implies algebraicity of the solutions. However, his argument is non-constructive. The purpose of this article is to explicitly construct a set of primes, for which it suffices to check the vanishing of the $p$-curvatures of a first order equation (2) – the cardinality of the set depending on a suitable measure of the ‘‘size’’ of $u(x)$ – to conclude that it has a non-zero algebraic solution. Our first main result is the following effective version of Kronecker’s Theorem.
@@
be a polynomial with leading coefficient $\Delta\coloneqq r_{n}>0$, and suppose that the maximal modulus of its complex roots is bounded by $B\in\mathbb{R}$. Let $\delta(\Delta)\coloneqq\prod_{p|\Delta}p^{1/(p-1)}$ and set $M\coloneqq\left\lceil 2.826\cdot\Delta^{3}\cdot\delta(\Delta)^{3}\right\rceil$ and $N\coloneqq\lceil 6.076\cdot BM\rceil$. Then $R(w)$ splits into linear factors in $\mathbb{Q}[w]$ if and only if its reduction modulo any prime number $p$ not dividing $\Delta$ and less than $\sigma\coloneqq(2M+1)N+2M$ splits into linear factors in $\mathbb{F}_{p}[w]$.
Our proof of this result makes the Chudnovskys’ proof of Kronecker’s result more explicit. The method of Hermite-Padй approximation inherently is effective in some sense, and in the conclusion of their article [CC85], the Chudnovsky brothers also claim that their results are effective. However most of their computations are sketched without many details, preventing us from directly deriving explicit bounds. Our contribution is to confirm their claim and to work out concrete bounds.
@@
and let $\Delta\coloneqq|\operatorname{res}(b(x),b^{\prime}(x))|=|r_{n}|$. With the notation as in Theorem 1.2, the equation $y^{\prime}(x)=u(x)y(x)$ has a non-zero algebraic solution if and only if its $p$-curvatures vanish for all primes not dividing $\Delta$ and smaller than $\sigma$.
The computational complexity of checking that sufficiently many $p$-curvatures vanish is $\tilde{O}(\Delta^{6}B)=\tilde{O}(H^{12n-6}n^{12n}3^{-3n})$ where the notation $\tilde{O}$ hides factors that are polynomial in $n$ and logarithmic in $H$.
@@
Moreover, we treat algorithmic aspects of the result and report on an implementation. We rely on a fast algorithm by Bostan and Schost [BS09] to quickly compute $p$-curvatures of order one equations. The algorithm presented in this text does not outperform other known algorithms for certifying that a differential equation of order one has algebraic solutions, however in generic cases the algorithm detects quickly the presence of a transcendental solution, especially when the degree and size of the coefficients explode. We expect that any significant improvement in our approach for certifying the algebraicity of solutions would come from new theoretical results rather than algorithmic optimization. Although theoretically the algorithm could treat the case of differential equations with algebraic non-rational coefficients, we did not optimize it to treat this case efficiently.
Our algorithm provides an instance of deciding algebraicity of D-finite functions using exclusively an arithmetic criterion. One could hope for results, similar in spirit, for differential equations of order one with algebraic coefficients, and for higher order equations in which the $p$-curvature conjecture is proven. The computation of $p$-curvatures for arbitrary order differential equations and deciding their nullity is possible and algorithms exist to perform these computations efficiently [BCS15, BCS16].
@@ Acknowledgments.
We are indebted to Alin Bostan, for his continuous support, for lots of insightful discussions and for many useful references. We thank Markus Reibnegger for his interest in the work, for pointing out crucial references, and for fruitful discussions.
@@ 2 Reformulations of the Problem
This section contains no new results. Its purpose is to describe the equivalence between the $p$-curvature conjecture for the order one equation (2) and Kronecker’s Theorem 1.1. The results can be found in one form or another in the literature [Hon81, CC85, Put96, BCR24].
We consider the equation (2) with $u(x)\in\mathbb{Q}(x)$. A nonzero solution $y(x)$ has the form $y(x)=\exp(\int u(x)\mathrm{d}x)$. This expression combines two operations that do not preserve rationality or algebraicity in different ways. A primitive of a rational function remains a rational function if and only if all its residues are zero, whereas a fraction $\frac{\alpha}{x-\beta}$ has a logarithm as a primitive. Contrarily, the exponential acts in such a way that for a nonzero algebraic function $f(x)$, the function $\exp(f(x))$ is transcendental. This can be seen as a particular case of a conjecture by Schanuel about the transcendence degree of a set of power series and their exponentials, that was proven by Ax [Ax71].
@@
A convenient way to compute the residues of the rational function $u(x)$ is provided by Rothstein-Trager resultants [Rot76, Rot77, Tra76]. We state their result here in a simplified, but for our purposes sufficient, form.
@@
Similar considerations hold in characteristic $p$. Moreover, by Cartier’s Lemma, the existence of a non-zero solution of an order one differential equation $y^{\prime}(x)=u(x)y(x)$ is equivalent to the vanishing of its $p$-curvature [^2] The existence of a non-zero polynomial solution in $\mathbb{F}_{p}[x]$, a non-zero algebraic solution in $\mathbb{F}_{p}[\![x]\!]$ or any non-zero solution in $\mathbb{F}_{p}[\![x]\!]$ are equivalent [Hon81, Lem. 1].. In the particular case of order one equations, the $p$-curvature is given by a one-dimensional matrix, whose entry we also call, by abuse of notation, the $p$-curvature. There is an explicit formula for this $p$-curvature in terms of the coefficient $u(x)$, that already implicitly appears in Jacobson’s work [Jac37]: it is given by $u(x)^{p}+u^{(p-1)}(x)$.
@@
Note that Honda actually proved more, namely that the $p$-curvature conjecture for order one equations is equivalent to Kronecker’s Theorem. We refer the reader for the other implication to Honda’s original article [Hon81].
Let us next investigate which prime numbers are exceptional in the sense that there is no solution of the reduction of the differential equation $y^{\prime}(x)=u(x)y(x)$ with $u(x)=a(x)/b(x)$ in $\mathbb{F}_{p}[\![x]\!],$ despite the existence of a solution in characteristic 0. In other words we are investigating for which prime numbers assertion (4) in Proposition 2.3 might be false, although (3) in Proposition 2.2 holds. Trivially, if a polynomial splits in $\mathbb{Z}[x]$ into linear factors, it splits into linear factors in $\mathbb{F}_{p}[x]$ not just for almost all prime numbers, but for all prime numbers. However, it might happen that the degree of $b(x)$ drops, when reduced modulo $p$, or that two distinct poles of $u(x)$ of order $1$ collapse to a pole of higher order, when reduced modulo $p$. Luckily, both of these exceptions can be easily read off the leading coefficient (up to sign) $\Delta=|\operatorname{res}(b(x),-b^{\prime}(x))|$ of the Rothstein-Trager resultant, as already noticed by van der Put, c.f. [Put96, Prop. 3.2(2)].
@@
Assume that $y^{\prime}(x)=u(x)y(x)$ with $u(x)=a(x)/b(x)$ has an algebraic solution. If $p\not\hskip 2.5pt\mid\Delta$ then the $p$-curvature of the equation vanishes.
@@
By Theorem 1.2, $R(w)$ factors into linear factors in $\mathbb{Q}[w]$ if and only if its reduction also splits completely in $\mathbb{F}_{p}[w]$ for all prime numbers $p$ up to $\sigma$, not dividing $\Delta$. By Proposition 2.3 this in turn is equivalent to the vanishing of the $p$-curvatures for all prime numbers not dividing $\Delta$ and smaller than $\sigma$.
@@
Let $K$ be a number field, let $R(w)\in K[w]$ with leading coefficient $\Delta\in K\backslash\{0\}$. If $R(w)$ splits completely over $K$ with all its roots in $\mathbb{Q}$, then $\frac{1}{\Delta}R(w)$ is a polynomial in $\mathbb{Q}[w]$.
@@ 3.1 Finding Rational Roots
A naïve approach to finding the rational roots of a polynomial is to factor it completely over $\mathbb{Q}$ and consider the factors of degree one. However this is more than what we are looking for and specific more efficient algorithms to find the rational roots of a polynomial exist. We refer the reader to [vzGG13, §14-16] and [Bos+17, §21] for details about such algorithms and an analysis of their complexity.

### v2

@@
Abstract
@@ 1 Introduction
Notably, in his paper from 1980 that popularized the study of D-finite series, i.e., power series that satisfiy a non-zero differential equation of the form (1), Stanley asked for an ‘‘algorithm suitable for computer implementation’’ for problem (P) [47].
Problem (A) was solved algorithmically in a seminal paper by Singer [46], relying on an algorithmic answer of Risch [39] in the case of first order differential equations with algebraic power series coefficients instead of polynomials – this instance of the problem is known as Abel’s problem – and on work conducted by Painlevй and Boulanger [38, 10]. However, this answer is not fully satisfactory, as the complexity of the algorithm involves exponential bounds, making it in general not suitable for implementation. Recent progress on Stanley’s question about problem (P) was made by Bostan, Salvy and Singer [8], giving (semi-) algorithms that rely on the minimization of differential operators annihilating a given D-finite series.
A different – arithmetic – approach to problem (A) was proposed by Grothendieck, in his $p$-curvature conjecture [23]. Given (1), we can pass to a matrix differential equation $Y^{\prime}(x)=A(x)Y(x)$ of order $1$, with $A(x)\in\mathbb{Q}(x)^{n\times n}$ being the companion matrix of (1). Let $p$ be a prime number. The $p$-curvature of the equation is defined as the $\mathbb{F}_{p}(x)$-linear map $\psi_{A}:\mathbb{F}_{p}(x)^{n}\to\mathbb{F}_{p}(x)^{n},Y(x)\mapsto(\partial-A_{p})^{p}Y(X)$, where $A_{p}(x)\in\mathbb{F}_{p}(x)^{n\times n}$ denotes the reduction of $A(x)$ modulo $p$. The $p$-curvature conjecture then states that (1) admits a basis of algebraic solutions if and only if its $p$-curvature vanishes for almost all, i.e., all but finitely many, prime numbers $p$. Cartier’s Lemma (as attributed by Katz [23]), asserts that the vanishing of the $p$-curvature is equivalent to the existence of a full basis of (algebraic) solutions of the reduction of the differential equation modulo $p$ in $\mathbb{F}_{p}[\![x]\!]$. Thus, the $p$-curvature conjecture can be seen as a local-global principle about the existence of algebraic solutions of a linear differential equation.
While in general wide open, the $p$-curvature conjecture is solved in many cases. Most notably, for ‘‘equations coming from geometry’’, i.e., suitable factors of Picard-Fuchs differential operators, a proof was given by Katz [23], and for hypergeometric differential equations an elementary proof is a by-product of the classification of algebraic hypergeometric functions [3, 15]. For first order equations
@@
with $u(x)\in\mathbb{Q}(x)$ the three problems (A), (E) and (P) are equivalent, as there is only a one dimensional solution space of the equation. Honda [18] provided an elementary proof of the $p$-curvature conjecture in this case, by showing that the problem is equivalent to a number theoretic result by Kronecker [25], which nowadays is often seen as a consequence of Chebotarev’s Density Theorem [53, 52, 49].
@@
For first order equations with algebraic power series coefficients, i.e., for the equation $y^{\prime}(x)=u(x)y(x)$ with $u(x)\in\overline{\mathbb{Q}}[\![x]\!]\cap\overline{\mathbb{Q}(x)}$, a proof of the $p$-curvature conjecture was given by the Chudnovsky brothers [11]. They use Hermite-Padй approximation to obtain their result. As a motivating example, they apply their methods to provide a new proof of Kronecker’s Theorem, not relying on Chebotarev’s Density Theorem, and in doing so giving also a new proof of Honda’s Theorem.
The $p$-curvature conjecture, as stated here, does not provide a way to decide whether (1) has a basis of algebraic solutions, as it translates the statement to a statement about almost all, in particular, a statement about infinitely many, prime numbers. In 1982 Katz extended the $p$-curvature conjecture to a conjecture about the Lie algebra of the differential Galois group of a differential equation [24]. From his work presented in [24, §9] it follows by a noetherianity argument that - in the cases in which his conjecture is true, among them order one equations with rational function coefficients - there exists a finite set of prime numbers, such that the vanishing of the $p$-curvatures modulo these primes implies algebraicity of the solutions. However, his argument is non-constructive. The purpose of this article is to explicitly construct a set of primes, for which it suffices to check the vanishing of the $p$-curvatures of a first order equation (2) – the cardinality of the set depending on a suitable measure of the ‘‘size’’ of $u(x)$ – to conclude that it has a non-zero algebraic solution. Our first main result is the following effective version of Kronecker’s Theorem.
@@
be a polynomial with leading coefficient $r_{n}>0$, let $\Delta_{R}\coloneqq\operatorname{res}_{w}(R_{0}(w),R_{0}^{\prime}(w))$, where $R_{0}(w)$ denotes the square-free part of $R(w)$, and suppose that the maximal modulus of its complex roots is bounded by $B\in\mathbb{R}$. Let $\delta(\Delta_{R})\coloneqq\prod_{p|\Delta_{R}}p^{1/(p-1)}$ and set $M\coloneqq\left\lceil 2.826\cdot r_{n}^{3}\cdot\delta(\Delta_{R})^{3}\right\rceil$ and $N\coloneqq\lceil 6.076\cdot BM\rceil$. Then $R(w)$ splits into linear factors in $\mathbb{Q}[w]$ if and only if its reduction modulo any prime number $p$ not dividing $r_{n}$ and less than $\sigma\coloneqq(2M+1)N+2M$ splits into linear factors in $\mathbb{F}_{p}[w]$.
Our proof of this result makes the Chudnovskys’ proof of Kronecker’s result more explicit. The method of Hermite-Padй approximation inherently is effective in some sense, and in the conclusion of their article [11], the Chudnovsky brothers also claim that their results are effective. However most of their computations are sketched without many details, preventing us from directly deriving explicit bounds. Our contribution is to confirm their claim and to work out concrete bounds.
@@
and let $\Delta_{b}\coloneqq|\operatorname{res}_{x}(b(x),b^{\prime}(x))|=|r_{n}|$. With the notation as in Theorem 1.2, the equation $y^{\prime}(x)=u(x)y(x)$ has a non-zero algebraic solution if and only if its $p$-curvatures vanish for all primes not dividing $\Delta_{b}$ and smaller than $\sigma$.
The computational complexity of checking that sufficiently many $p$-curvatures vanish is $\tilde{O}(\Delta_{b}^{6}B)=\tilde{O}(H^{12n-6}n^{12n}3^{-3n})$ where the notation $\tilde{O}$ hides factors that are polynomial in $n$ and logarithmic in $H$.
@@
Moreover, we treat algorithmic aspects of the result and report on an implementation. We rely on a fast algorithm by Bostan and Schost [9] to quickly compute $p$-curvatures of order one equations. The algorithm presented in this text does not outperform other known algorithms for certifying that a differential equation of order one has algebraic solutions, however in generic cases the algorithm detects quickly the presence of a transcendental solution, especially when the degree and size of the coefficients explode. We expect that any significant improvement in our approach for certifying the algebraicity of solutions would come from new theoretical results rather than algorithmic optimization. Although theoretically the algorithm could treat the case of differential equations with algebraic non-rational coefficients, we did not optimize it to treat this case efficiently.
Our algorithm provides an instance of deciding algebraicity of D-finite functions using exclusively an arithmetic criterion. One could hope for results, similar in spirit, for differential equations of order one with algebraic coefficients, and for higher order equations in which the $p$-curvature conjecture is proven. The computation of $p$-curvatures for arbitrary order differential equations and deciding their nullity is possible and algorithms exist to perform these computations efficiently [5, 6].
@@ Acknowledgments.
We are indebted to Alin Bostan, for his continuous support, for lots of insightful discussions and for many useful references, and to Raphaël Pagиs for pointing out a serious flaw in Lemma 4.4 in an earlier version of this text, and for his help with repairing it. Moreover, we thank Markus Reibnegger for his interest in the work, for pointing out crucial references, and for fruitful discussions.
@@ 2 Reformulations of the Problem
This section contains no new results. Its purpose is to describe the equivalence between the $p$-curvature conjecture for the order one equation (2) and Kronecker’s Theorem 1.1. The results can be found in one form or another in the literature [18, 11, 51, 4].
We consider the equation (2) with $u(x)\in\mathbb{Q}(x)$. A nonzero solution $y(x)$ has the form $y(x)=\exp(\int u(x)\mathrm{d}x)$. This expression combines two operations that do not preserve rationality or algebraicity in different ways. A primitive of a rational function remains a rational function if and only if all its residues are zero, whereas a fraction $\frac{\alpha}{x-\beta}$ has a logarithm as a primitive. Contrarily, the exponential acts in such a way that for a nonzero algebraic function $f(x)$, the function $\exp(f(x))$ is transcendental. This can be seen as a particular case of a conjecture by Schanuel about the transcendence degree of a set of power series and their exponentials, that was proven by Ax [2].
@@
A convenient way to compute the residues of the rational function $u(x)$ is provided by Rothstein-Trager resultants [42, 43, 48]. We state their result here in a simplified, but for our purposes sufficient, form.
@@
Similar considerations hold in characteristic $p$. Moreover, by Cartier’s Lemma, the existence of a non-zero solution of an order one differential equation $y^{\prime}(x)=u(x)y(x)$ is equivalent to the vanishing of its $p$-curvature [^2] The existence of a non-zero polynomial solution in $\mathbb{F}_{p}[x]$, a non-zero algebraic solution in $\mathbb{F}_{p}[\![x]\!]$ or any non-zero solution in $\mathbb{F}_{p}[\![x]\!]$ are equivalent [18, Lem. 1].. In the particular case of order one equations, the $p$-curvature is given by a one-dimensional matrix, whose entry we also call, by abuse of notation, the $p$-curvature. There is an explicit formula for this $p$-curvature in terms of the coefficient $u(x)$, that already implicitly appears in Jacobson’s work [19]: it is given by $u(x)^{p}+u^{(p-1)}(x)$.
@@
Note that Honda actually proved more, namely that the $p$-curvature conjecture for order one equations is equivalent to Kronecker’s Theorem. We refer the reader for the other implication to Honda’s original article [18].
Let us next investigate which prime numbers are exceptional in the sense that there is no solution of the reduction of the differential equation $y^{\prime}(x)=u(x)y(x)$ with $u(x)=a(x)/b(x)$ in $\mathbb{F}_{p}[\![x]\!],$ despite the existence of a solution in characteristic 0. In other words we are investigating for which prime numbers assertion (4) in Proposition 2.3 might be false, although (3) in Proposition 2.2 holds. Trivially, if a polynomial splits in $\mathbb{Z}[x]$ into linear factors, it splits into linear factors in $\mathbb{F}_{p}[x]$ not just for almost all prime numbers, but for all prime numbers. However, it might happen that the degree of $b(x)$ drops, when reduced modulo $p$, or that two distinct poles of $u(x)$ of order $1$ collapse to a pole of higher order, when reduced modulo $p$. Luckily, both of these exceptions can be easily read off the leading coefficient (up to sign) $\Delta_{b}=|\operatorname{res}(b(x),-b^{\prime}(x))|$ of the Rothstein-Trager resultant, as already noticed by van der Put, c.f. [51, Prop. 3.2(2)].
@@
Assume that $y^{\prime}(x)=u(x)y(x)$ with $u(x)=a(x)/b(x)$ has an algebraic solution. If $p\not\hskip 2.5pt\mid\Delta_{b}$ then the $p$-curvature of the equation vanishes.
@@
By Theorem 1.2, $R(w)$ factors into linear factors in $\mathbb{Q}[w]$ if and only if its reduction also splits completely in $\mathbb{F}_{p}[w]$ for all prime numbers $p$ up to $\sigma$, not dividing $\Delta_{b}$. By Proposition 2.3 this in turn is equivalent to the vanishing of the $p$-curvatures for all prime numbers not dividing $\Delta_{b}$ and smaller than $\sigma$.
@@
Let $K$ be a number field, let $R(w)\in K[w]$ with leading coefficient $r_{n}\in K\backslash\{0\}$. If $R(w)$ splits completely over $K$ with all its roots in $\mathbb{Q}$, then $\frac{1}{r_{n}}R(w)$ is a polynomial in $\mathbb{Q}[w]$.
@@ 3.1 Finding Rational Roots
A naïve approach to finding the rational roots of a polynomial is to factor it completely over $\mathbb{Q}$ and consider the factors of degree one. However this is more than what we are looking for and specific more efficient algorithms to find the rational roots of a polynomial exist. We refer the reader to [13, §14-16] and [7, §21] for details about such algorithms and an analysis of their complexity.